OpenStudy (anonymous):

What is the equation of a circle with center (6,4) that passes through the coordinate (2,1)?

OpenStudy (mathmale):

Would you look up and type out here the general equation of a circle with center at (h,k) and radius r, please? Then we'll get to work on your math problem.

OpenStudy (anonymous):

circle: (x - h)2 + (y - k)2 = r2 radius: is the r2 in that equation.

OpenStudy (mathmale):

OK. Good. You are told that the center of your circle is (6,4). Modify your equation accordingly. Type out the result here, please.

OpenStudy (anonymous):

Im confused

OpenStudy (anonymous):

I don't know how to set up the equation.

OpenStudy (mathmale):

You were correct in typing out (x - h)2 + (y - k)2 = r2, which in more formal notation is\[(x - h)^2 + (y - k)^2 = r^2\] You told me that the center of your particular circle is (6,4). That means your h=6 and your k=4. I'm asking you to substitute those values of h and k into your \((x - h)^2 + (y - k)^2 = r^2\) now.

OpenStudy (anonymous):

i figured it out. the 6,4 is the k,h so i plugged it in and the other coordinates are x,y i plugged them in.